# Estimators for multivariate allometric regression model

### v2

@@ 3 Estimators based on weighted matrices of $\bm{S}_{R}$ and $\bm{S}_{E}$
$=\frac{8\{q(1-2w)+(n-1)w^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w)^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w\}(\lambda_{1}-\lambda_{2})+(1-w)\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}},$
@@
$\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(w)-\bm{\gamma}_{1}\|\stackrel{{\scriptstyle\mathrm{P}}}{{\to}}0$ ( 3.2)
for any value of parameters satisfying underlying assumptions, where $\stackrel{{\scriptstyle\mathrm{P}}}{{\to}}$ denotes convergence in probability. Note that (3.2) is equivalent to $\mathsf{E}[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(w)-\bm{\gamma}_{1}\|^{2}]\to 0$, because $0\leq\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(w)-\bm{\gamma}_{1}\|\leq\sqrt{2}$. Clearly, whether $\bm{\alpha}=\bm{0}_{q}$ or not is crucial for $\hat{\bm{\gamma}}_{1}(w)$ with $w\in[0,1)$, so, in some settings, consistency will be discussed under the assumption that $\bm{\alpha}\neq\bm{0}_{q}$.
@@
We use Lemma 3.1 to bound $\mathsf{E}[\|\bm{S}(w)-\mathsf{E}[\bm{S}(w)]\|^{2}_{\rm op}]$ for the purpose of getting a specific bound that will be used to determine the value of $w$ below. Evaluating $\mathsf{E}[\|\bm{S}(w)-\mathsf{E}[\bm{S}(w)]\|_{\rm op}]$ directly may provide a better bound. When $w=1$ or when $\bm{\alpha}=\bm{0}_{q},w=1/2$, results such as Theorem 4 of Koltchinskii and Lounici (2017) or Theorem 4.7.1 of Vershynin (2018) can be used. For example, when $\bm{\alpha}=\bm{0}_{q}$, it holds that
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}^{(T)}-\bm{\gamma}_{1}\|^{2}\right]\leq\mathsf{E}\left[\sqrt{2}\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}^{(T)}-\bm{\gamma}_{1}\|\right]\leq\frac{C}{\lambda_{1}-\lambda_{2}}\max\left\{\sqrt{\frac{\lambda_{1}{\rm tr}(\bm{\Sigma})}{n-1}},\frac{{\rm tr}(\bm{\Sigma})}{n-1}\right\},$
@@
$w^{*}=\frac{q\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}(\lambda_{1}-\lambda_{2})+2(\lambda_{1}+{\rm tr}(\bm{\Sigma}))(\lambda_{1}-\lambda_{2})\|\bm{X}\bm{\alpha}\|^{2}}{2q\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}(\lambda_{1}-\lambda_{2})+2(\lambda_{1}+{\rm tr}(\bm{\Sigma}))(\lambda_{1}-\lambda_{2})\|\bm{X}\bm{\alpha}\|^{2}+\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}\|\bm{X}\bm{\alpha}\|^{2}}.$ ( 3.3)
@@
For simplicity, we let
$a={\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2},\ \ \ b=\lambda_{1}+{\rm tr}(\bm{\Sigma}),\ \ \ c=\|\bm{X}\bm{\alpha}\|^{2},\ \ \ d=\lambda_{1}-\lambda_{2}.$
Then, (3.1) is rewritten as
$8\frac{\{q(1-2w)+(n-1)w^{2}\}a+2(1-w)^{2}bc}{\left[\{q+(n-1-2q)w\}d+(1-w)c\right]^{2}},$
@@
$f(w)=\frac{\{q(1-2w)+(n-1)w^{2}\}a+2(1-w)^{2}bc}{2\left[\{q+(n-1-2q)w\}d+(1-w)c\right]^{2}}.$
Since $\{q+(n-1-2q)w\}d>0$ for $0\leq w\leq 1$ due to $q+(n-1-2q)w=q(1-w)+(n-1-q)w>0$ and $d=\lambda_{1}-\lambda_{2}>0$, the function $f(w)$ is continuous on $0\leq w\leq 1$. Differentiating $f(w)$ with respect to $w$, we have
$\frac{\mathrm{d}f(w)}{\mathrm{d}w}$ =$=\frac{\{-2q+2(n-1)w\}a-4(1-w)bc}{2\left[\{q+(n-1-2q)w\}d+(1-w)c\right]^{2}}-\frac{\left[\{q(1-2w)+(n-1)w^{2}\}a+2(1-w)^{2}bc\right]\left\{(n-1-2q)d-c\right\}}{\left[\{q+(n-1-2q)w\}d+(1-w)c\right]^{3}}$
=$=\frac{\left\{(2adq+ac+2bcd)w-(adq+2bcd)\right\}(n-1-q)}{\left[\{q+(n-1-2q)w\}d+(1-w)c\right]^{3}},$
@@
Hence, the value of $w$ minimizing the function $f(w)$ is given by
$w^{*}=\frac{adq+2bcd}{2adq+2bcd+ac}.$
Substituting $a,b,c,d$, we have (3.3). This completes the proof. ∎
@@
$(\lambda_{1}+{\rm tr}(\bm{\Sigma}))(\lambda_{1}-\lambda_{2})<\lambda_{1}^{2}+\lambda_{1}{\rm tr}(\bm{\Sigma})<{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}$
@@
$w^{*}$ =$=\frac{\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}(\lambda_{1}-\lambda_{2})q+2(\lambda_{1}+{\rm tr}(\bm{\Sigma}))(\lambda_{1}-\lambda_{2})\|\bm{X}\bm{\alpha}\|^{2}}{2\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}(\lambda_{1}-\lambda_{2})q+2(\lambda_{1}+{\rm tr}(\bm{\Sigma}))(\lambda_{1}-\lambda_{2})\|\bm{X}\bm{\alpha}\|^{2}+\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}\|\bm{X}\bm{\alpha}\|^{2}}$
<$<\frac{\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}(\lambda_{1}-\lambda_{2})q+2(\lambda_{1}+{\rm tr}(\bm{\Sigma}))(\lambda_{1}-\lambda_{2})\|\bm{X}\bm{\alpha}\|^{2}}{2\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}(\lambda_{1}-\lambda_{2})q+3(\lambda_{1}+{\rm tr}(\bm{\Sigma}))(\lambda_{1}-\lambda_{2})\|\bm{X}\bm{\alpha}\|^{2}}\leq\frac{2}{3}.$
@@ 4 Consistency under weak identifiability regimes
In this section, we discuss the case of $n\to\infty$ with fixed $p,q$. Assume that $n^{-1}\bm{X}^{\prime}\bm{X}$ converges to a positive-definite matrix as $n\to\infty$. Let $\bm{\alpha}$ be fixed. Then, $\|\bm{X}\bm{\alpha}\|^{2}=\Theta(n)$ holds for $\bm{\alpha}\neq\bm{0}_{q}$, and $\|\bm{X}\bm{\alpha}\|^{2}=0$ holds for $\bm{\alpha}=\bm{0}_{q}$, where $a_{n}=\Theta(n^{u})$ for a sequence $\{a_{n}\}_{n\in\mathbb{N}}$ means that lim $\liminf(a_{n}/n^{u})>0$ and lim $\limsup(a_{n}/n^{u})<\infty$ as $n\to\infty$. If $\bm{\Sigma}$ is fixed (with $\lambda_{1}-\lambda_{2}>0$), then it is obvious that $\hat{\bm{\gamma}}_{1}(w)$ is always a consistent estimator of $\bm{\gamma}_{1}$ up to sign for all $w\in[0,1]$, because the upper bound (3.1) for the mean squared error of $\hat{\bm{\gamma}}_{1}(w)$ converges to 00 as follows:
@@
$\leq\frac{8\{q(1-2w)+(n-1)w^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w)^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w\}(\lambda_{1}-\lambda_{2})+(1-w)\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}$
$=\frac{8\{q(1-w)^{2}+(n-1-q)w^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w)^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q(1-w)+(n-1-q)w\}(\lambda_{1}-\lambda_{2})+(1-w)\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}$
$=\frac{\Theta(n)}{\left(\Theta(n)\right)^{2}}=\Theta\left(n^{-1}\right).$
Hence, in this section, we focus on the case of $\lambda_{1}-\lambda_{2}\to 0$ as $n\to\infty$. Such regimes are called weak identifiability regimes in Paindaveine et al. (2020) and Paindaveine and Verdebout (2023) in which some specific structures for $\bm{\Sigma}$ are assumed. These regimes correspond to the cases where the eigenvector of $\bm{\Sigma}$ corresponding to the largest eigenvalue is not identifiable in the limit. More specifically, we consider asymptotic regimes $n\to\infty$ under the following assumption.
@@
[…]
@@
First, we discuss the case of $\bm{\alpha}\neq\bm{0}_{q}$, that is, $\|\bm{X}\bm{\alpha}\|^{2}=\Theta(n)$. If $10<w<$ 1, then
@@
$\leq\frac{8\{q(1-2w)+(n-1)w^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w)^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w\}(\lambda_{1}-\lambda_{2})+(1-w)\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}$
@@
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(0)-\bm{\gamma}_{1}\|^{2}\right]$ $\leq\frac{8q\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left\{q(\lambda_{1}-\lambda_{2})+\|\bm{X}\bm{\alpha}\|^{2}\right\}^{2}}$
$=\frac{\Theta(1)+\Theta(n)}{\left(\Theta(n^{-\eta})+\Theta(n)\right)^{2}}=\frac{\Theta(n)}{\left(\Theta(n)\right)^{2}}=\Theta\left(n^{-1}\right).$
@@
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(1)-\bm{\gamma}_{1}\|^{2}\right]$ $\leq\frac{8(n-1-q)\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}}{\left\{(n-1-q)(\lambda_{1}-\lambda_{2})\right\}^{2}}$
$=\frac{\Theta(n)}{\left(\Theta(n^{1-\eta})\right)^{2}}=\frac{\Theta(n)}{\Theta(n^{2-2\eta})}=\Theta\left(n^{2\eta-1}\right).$
@@
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(w)-\bm{\gamma}_{1}\|^{2}\right]$ $\leq\frac{8\{q(1-2w)+(n-1)w^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}}{\left[\{q+(n-1-2q)w\}(\lambda_{1}-\lambda_{2})\right]^{2}}$
$=\frac{\Theta(n)}{\left(\Theta(n^{1-\eta})\right)^{2}}=\frac{\Theta(n)}{\Theta(n^{2-2\eta})}=\Theta\left(n^{2\eta-1}\right).$
@@
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(0)-\bm{\gamma}_{1}\|^{2}\right]\leq\frac{8q\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}}{\left\{q(\lambda_{1}-\lambda_{2})\right\}^{2}}=\frac{\Theta(1)}{\left(\Theta(n^{-\eta})\right)^{2}}=\Theta\left(n^{2\eta}\right).$
@@
From Remarks 1 and 5, when $\bm{\alpha}=\bm{0}_{q}$, if $\lambda_{1}-\lambda_{2}\to 0$ and $n^{1/2}(\lambda_{1}-\lambda_{2})\to\infty$, then the leading principal eigenvector $\hat{\bm{\gamma}}_{1}^{(T)}$ is consistent. It is consistent with a result in Paindaveine et al. (2020, Lemma 2.3). Also from Paindaveine et al. (2020, Lemma 2.3), if $\limsup_{n\to\infty}n^{1/2}(\lambda_{1}-\lambda_{2})<\infty$ then $\hat{\bm{\gamma}}_{1}^{(T)}$ is no more consistent, because their setting is a special case of ours.
@@
We notice that $w^{*}=1/2$ under $\bm{\alpha}=\bm{0}_{q}$ and $w^{*}=\Theta(n^{-\eta})$ under $\bm{\alpha}\neq\bm{0}_{q}$. For $\bm{\alpha}=\bm{0}_{q}$, since $w^{*}=1/2$, we see from Theorem 4.1 that $\eta<1/2$ is a sufficient condition for (4.1). For $\bm{\alpha}\neq\bm{0}_{q}$,
@@
$\leq\frac{8\{q(1-w^{*})^{2}+(n-1-q)(w^{*})^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w^{*})^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q(1-w^{*})+(n-1-q)w^{*}\}(\lambda_{1}-\lambda_{2})+(1-w^{*})\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}$

### v4

@@ 3 Estimators based on weighted matrices of $\bm{S}_{R}$ and $\bm{S}_{E}$
$=\frac{8a\{q(1-2w)+(n-1)w^{2}\}+16bc(1-w)^{2}}{\left[d\{q+(n-1-2q)w\}+c(1-w)\right]^{2}},$
@@
$\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(w)-\bm{\gamma}_{1}\|\stackrel{{\scriptstyle\mathrm{P}}}{{\to}}0$ ( 3.3)
for any value of parameters satisfying underlying assumptions, where $\stackrel{{\scriptstyle\mathrm{P}}}{{\to}}$ denotes convergence in probability. Note that (3.3) is equivalent to
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(w)-\bm{\gamma}_{1}\|^{2}\right]\to 0,$
because
$0\leq\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(w)-\bm{\gamma}_{1}\|\leq\sqrt{2}.$
Whether $\bm{\alpha}=\bm{0}_{q}$ or not is crucial for $\hat{\bm{\gamma}}_{1}(w)$ with $w\in[0,1)$, so, in some settings, consistency will be discussed under the assumption that $\bm{\alpha}\neq\bm{0}_{q}$.
@@
We use Lemma 3.1 to bound $\mathsf{E}[\|\bm{S}(w)-\mathsf{E}[\bm{S}(w)]\|^{2}_{\rm op}]$ for the purpose of getting a specific bound that will be used to determine the value of $w$ below. Evaluating $\mathsf{E}[\|\bm{S}(w)-\mathsf{E}[\bm{S}(w)]\|_{\rm op}]$ directly may provide a better bound. When $w=1$ or when $\bm{\alpha}=\bm{0}_{q},w=1/2$, results such as Koltchinskii and Lounici (2017, Theorem 4) or Vershynin (2018, Theorem 4.7.1) can be used. For example, when $\bm{\alpha}=\bm{0}_{q}$, it holds that
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}^{(T)}-\bm{\gamma}_{1}\|^{2}\right]$
$\leq\mathsf{E}\left[\sqrt{2}\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}^{(T)}-\bm{\gamma}_{1}\|\right]\leq\frac{C}{\lambda_{1}-\lambda_{2}}\max\left\{\sqrt{\frac{\lambda_{1}{\rm tr}(\bm{\Sigma})}{n-1}},\frac{{\rm tr}(\bm{\Sigma})}{n-1}\right\},$
@@
$w^{*}=\frac{adq+2bcd}{2adq+2bcd+ac},$ ( 3.4)
where $a$, $b$, $c$, and $d$ are given in (3.2).
@@
The right-hand side of (3.1) is rewritten as
$8\frac{a\{q(1-2w)+(n-1)w^{2}\}+2bc(1-w)^{2}}{\left[d\{q+(n-1-2q)w\}+c(1-w)\right]^{2}},$ ( 3.5)
@@
$f(w)=\frac{a\{q(1-2w)+(n-1)w^{2}\}+2bc(1-w)^{2}}{2\left[d\{q+(n-1-2q)w\}+c(1-w)\right]^{2}}.$
Since $d\{q+(n-1-2q)w\}>0$ for $0\leq w\leq 1$ due to $q+(n-1-2q)w=q(1-w)+(n-1-q)w>0$ and $d=\lambda_{1}-\lambda_{2}>0$, the function $f(w)$ is continuous on $0\leq w\leq 1$. Differentiating $f(w)$ with respect to $w$, we have
$\frac{\mathrm{d}f(w)}{\mathrm{d}w}$ $=\frac{a\{-2q+2(n-1)w\}-4bc(1-w)}{2\left[d\{q+(n-1-2q)w\}+c(1-w)\right]^{2}}$
$\qquad-\frac{\left[a\{q(1-2w)+(n-1)w^{2}\}+2bc(1-w)^{2}\right]\left\{d(n-1-2q)-c\right\}}{\left[d\{q+(n-1-2q)w\}+c(1-w)\right]^{3}}$
=$=\frac{\left\{(2adq+ac+2bcd)w-(adq+2bcd)\right\}(n-1-q)}{\left[d\{q+(n-1-2q)w\}+c(1-w)\right]^{3}},$
@@
Hence, the value of $w$ minimizing the function $f(w)$ is given by (3.4). This completes the proof. ∎
@@
$bd=(\lambda_{1}+{\rm tr}(\bm{\Sigma}))(\lambda_{1}-\lambda_{2})<\lambda_{1}^{2}+\lambda_{1}{\rm tr}(\bm{\Sigma})<{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}=a$
@@
$w^{*}=\frac{adq+2bcd}{2adq+2bcd+ac}<\frac{adq+2bcd}{2adq+3bcd}\leq\frac{2}{3}.$
@@ 4 Consistency under weak identifiability regimes
In this section, we discuss the case of $n\to\infty$ with fixed $p,q$. Assume that $n^{-1}\bm{X}^{\prime}\bm{X}$ converges to a positive-definite matrix as $n\to\infty$. Let $\bm{\alpha}$ be fixed. Then, $\|\bm{X}\bm{\alpha}\|^{2}=\Theta(n)$ holds for $\bm{\alpha}\neq\bm{0}_{q}$, and $\|\bm{X}\bm{\alpha}\|^{2}=0$ holds for $\bm{\alpha}=\bm{0}_{q}$, where $a_{n}=\Theta(n^{u})$ for a sequence $\{a_{n}\}_{n\in\mathbb{N}}$ means that lim $\liminf(a_{n}/n^{u})>0$ and lim $\limsup(a_{n}/n^{u})<\infty$ as $n\to\infty$. If $\bm{\Sigma}$ is fixed (with $\lambda_{1}-\lambda_{2}>0$), then $\hat{\bm{\gamma}}_{1}(w)$ is always a consistent estimator of $\bm{\gamma}_{1}$ up to sign for all $w\in(0,1]$, because the upper bound for the mean squared error of $\hat{\bm{\gamma}}_{1}(w)$ in (3.1) converges to 00 as follows:
@@
$\leq\frac{8a\{q(1-2w)+(n-1)w^{2}\}+16bc(1-w)^{2}}{\left[d\{q+(n-1-2q)w\}+c(1-w)\right]^{2}}$
$=\frac{8a\{q(1-w)^{2}+(n-1-q)w^{2}\}+16bc(1-w)^{2}}{\left[d\{q(1-w)+(n-1-q)w\}+c(1-w)\right]^{2}}=\frac{\Theta(n)}{\left(\Theta(n)\right)^{2}}=\Theta\left(n^{-1}\right),$
where the notation (3.2) is used. When $\bm{\alpha}\neq\bm{0}_{q}$, so is $\hat{\bm{\gamma}}_{1}(0)$. Hence, in this section, we focus on the case of $\lambda_{1}-\lambda_{2}\to 0$ as $n\to\infty$. Such regimes are called weak identifiability regimes in Paindaveine et al. (2020) and Paindaveine and Verdebout (2023) in which some specific structures for $\bm{\Sigma}$ are assumed. These regimes correspond to the cases where the eigenvector of $\bm{\Sigma}$ corresponding to the largest eigenvalue is not identifiable in the limit. More specifically, we consider asymptotic regimes $n\to\infty$ under the following assumption.
@@
$\begin{array}[]{ll}\bm{\alpha}\neq\bm{0}_{q}&\mbox{for}\ w=0,\\ \bm{\alpha}\neq\bm{0}_{q}\ \mbox{or}\ \eta<1/2&\mbox{for}\ 0<w<1,\\ \eta<1/2&\mbox{for}\ w=1.\end{array}$
@@
In this proof, we use the notation (3.2) for simplicity. We notice that $a=\Theta(1)$, $b=\Theta(1)$, $d=\Theta(n^{-\eta})$, and that $c=0$ under $\bm{\alpha}=\bm{0}_{q}$ and $c=\Theta(n)$ under $\bm{\alpha}\neq\bm{0}_{q}$. First, we discuss the case of $\bm{\alpha}\neq\bm{0}_{q}$, that is, $\|\bm{X}\bm{\alpha}\|^{2}=\Theta(n)$. If $10<w<$ 1, then
@@
$\leq\frac{8a\{q(1-2w)+(n-1)w^{2}\}+16bc(1-w)^{2}}{\left[d\{q+(n-1-2q)w\}+c(1-w)\right]^{2}}$
@@
$\leq\frac{8aq+16bc}{(dq+c)^{2}}=\frac{\Theta(1)+\Theta(n)}{\left(\Theta(n^{-\eta})+\Theta(n)\right)^{2}}=\frac{\Theta(n)}{\left(\Theta(n)\right)^{2}}=\Theta\left(n^{-1}\right).$
@@
$\leq\frac{8a(n-1-q)}{d^{2}(n-1-q)^{2}}=\frac{\Theta(n)}{\left(\Theta(n^{1-\eta})\right)^{2}}=\frac{\Theta(n)}{\Theta(n^{2-2\eta})}=\Theta\left(n^{2\eta-1}\right).$
@@
$\leq\frac{8a\{q(1-2w)+(n-1)w^{2}\}}{d^{2}\{q+(n-1-2q)w\}^{2}}=\frac{\Theta(n)}{\left(\Theta(n^{1-\eta})\right)^{2}}=\frac{\Theta(n)}{\Theta(n^{2-2\eta})}=\Theta\left(n^{2\eta-1}\right).$
@@
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(0)-\bm{\gamma}_{1}\|^{2}\right]\leq\frac{8aq}{d^{2}q^{2}}=\frac{\Theta(1)}{\left(\Theta(n^{-\eta})\right)^{2}}=\Theta\left(n^{2\eta}\right).$
@@
From Remarks 2 and 6, when $\bm{\alpha}=\bm{0}_{q}$, if $\lambda_{1}-\lambda_{2}\to 0$ and $n^{1/2}(\lambda_{1}-\lambda_{2})\to\infty$, then consistency up to sign of the leading principal eigenvector $\hat{\bm{\gamma}}_{1}^{(T)}$ holds. It is consistent with a result in Paindaveine et al. (2020, Lemma 2.3). Also from Paindaveine et al. (2020, Lemma 2.3), if $\limsup_{n\to\infty}n^{1/2}(\lambda_{1}-\lambda_{2})<\infty$ then $\hat{\bm{\gamma}}_{1}^{(T)}$ is no more consistent, because their setting is a special case of ours.
@@
In this proof, we use the notation (3.2) for simplicity, and remark that $w^{*}=1/2$ under $\bm{\alpha}=\bm{0}_{q}$ and that $w^{*}=\Theta(n^{-\eta})$ under $\bm{\alpha}\neq\bm{0}_{q}$. For $\bm{\alpha}=\bm{0}_{q}$, since $w^{*}=1/2$, we see from Theorem 4.1 that $\eta<1/2$ is a sufficient condition for (4.1). For $\bm{\alpha}\neq\bm{0}_{q}$,
@@
$\leq\frac{8a\{q(1-w^{*})^{2}+(n-1-q)(w^{*})^{2}\}+16bc(1-w^{*})^{2}}{\left[d\{q(1-w^{*})+(n-1-q)w^{*}\}+c(1-w^{*})\right]^{2}}$